x*e^-x +1=0. find no.of roots in {-1,-1/2}

Grade:upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
f(x)=x*e^(-x)+1 e=2.73
f(-1)=-e+1 <0
f(-1/2)=1/2e^(-1/2)+1 >0

f(x) monotonic function , imply there exitst a one root of function.

sher mohammad
b.tech, iit delhi